Hi, I'm trying to learn how to use the Arrhenius equation and I am very confused, can someone help? Thank you!

The activation energy for the gas phase decomposition of cyclobutane is 262 kJ/mol.

(CH2)4ebab9d75-6e9a-4c80-9bcd-b491b668f097.gif2 C2H4


The rate constant for this reaction is 2.16×10-4 s-1 at 711 K. What is the rate constant at 749 K?

Use the Arrhenius equation to determine Ea.

Close Problem

For the gas phase decomposition of cyclobutane, the rate constant is 8.84×10-4 s-1 at 734 K and 5.91×10-3 s-1 at 768 K.

  • (CH2)4ebab9d75-6e9a-4c80-9bcd-b491b668f097.gif2 C2H4


Calculate the activation energy for this reaction.  kJ/mol

Use Arrhenius plot information to determine Ea.

Close Problem

The rate of the following reaction was measured at different temperatures and rate constants were determined.

  • CH3CHClCH2Clebab9d75-6e9a-4c80-9bcd-b491b668f097.gifCH3CH=CHCl + HCl

Which of the following should give a linear plot (temperatures in K)?

On the x-axis: _____Tln T1/T   On the y-axis: _____kln k1/k



When the appropriate plot is made, the resulting least-squares line has a slope of -2.77×104 K and a y-intercept of 31.8.

Use this information to determine the activation energy for the reaction.

Ea =  kJ/mol
Answer #1

Arrhenius equation gives dependence of the rate constant on temperature.

Arrhenius equation :

k=Ae-Ea/RT

k= rate constant

A= pre-exponential factor

Ea =Activation energy

T= temperature

R= universal gas constant

At temperature T1 the rate of reaction becomes k1 and at temperature T2 the rate of reaction becomes k2.

So, k1=Ae-Ea/RT1

lnk1 = lnA - Ea/RT1...................(1)

and k2=Ae-Ea/RT2

lnk2 = lnA - Ea/RT2...................(2)

subtracting (2) from (1)

lnk1 - lnk2 = - Ea/RT1 + Ea/RT2

ln(k1 / k2)= Ea/R(1/ T2 -1/T1 )

ln(k1 / k2)=Ea/RT1T2 (T1-T2)

  • The rate constant for gas phase decomposition of cyclobutane at 749K is

ln(2.16×10-4 /k2)= 262*103J* (711K-749K) / 8.314 J mol-1 K-1 *711K*749K

-8.4402-lnk2 = 0.05917*(-38)

lnk2 = -8.4402+2.2486

lnk2 =-6.1915

k2 = 2.0467*10-3

  • The activation energy for the reaction is

ln(8.84×10-4 /5.91×10-3)=Ea (734K-768K) / 8.314 J mol-1 K-1 *734K*768K

-1.8998= Ea (-34) / 4.6867*106

Ea = 272.579 KJ

  • lnk1 = lnA - Ea/RT1

This equation is like y=mx+c where m= slope and c=intercept

So, the plot of lnk1 vs 1/T1 will be linear in nature where y=lnk1 and x = 1/T1

slope= - Ea/R and intercept = lnA

- Ea/R =-2.77×104K

Ea= 2.77×104K* 8.314 J mol-1 K-1

Ea= 230.2978 KJ mol-1